OpenStudy (anonymous):

tell between which 2 consecutive integers the square root of 67 lie

OpenStudy (anonymous):

Well lets start off with \[\sqrt{67}\] what do we get?

OpenStudy (tkhunny):

\(1^{2} = 1\) \(2^{2} = 4\) \(3^{2} = 9\) \(4^{2} = 16\) \(5^{2} = 25\) \(6^{2} = 26\) \(7^{2} = 49\) \(8^{2} = 64\) FOUND IT! It's in here, somewhere. No need to calculate it's value. Just answer the question. \(9^{2} = 81\)

OpenStudy (anonymous):

I think I got it. The square root of 67 lies between 8 and 9?

jhonyy9 (jhonyy9):

@tkhunny 6^2 =36 - not 26 - sorry for my attention

OpenStudy (tkhunny):

I'm delighted that you actually read it.